### v1

# Beyond the Leaderboard: Rethinking Medical Benchmarks for Large Language Models

Abstract
Large language models (LLMs) show significant potential in healthcare, prompting numerous benchmarks to evaluate their capabilities. However, concerns persist regarding the reliability of these benchmarks, which often lack clinical fidelity, robust data management, and safety-oriented evaluation metrics. To address these shortcomings, we introduce MedCheck, the first lifecycle-oriented assessment framework specifically designed for medical benchmarks. Our framework deconstructs a benchmark’s development into five continuous stages, from design to governance, and provides a comprehensive checklist of 46 medically-tailored criteria. Using MedCheck, we conducted an in-depth empirical evaluation of 53 medical LLM benchmarks. Our analysis uncovers widespread, systemic issues, including a profound disconnect from clinical practice, a crisis of data integrity due to unmitigated contamination risks, and a systematic neglect of safety-critical evaluation dimensions like model robustness and uncertainty awareness. Based on these findings, MedCheck serves as both a diagnostic tool for existing benchmarks and an actionable guideline to foster a more standardized, reliable, and transparent approach to evaluating AI in healthcare.
Beyond the Leaderboard: Rethinking Medical Benchmarks for Large Language Models
Zizhan Ma1  Wenxuan Wang2  Guo Yu3  Yiu-Fai Cheung 1 Meidan Ding3  Jie Liu4  Wenting Chen4  Linlin Shen3 1 The Chinese University of Hong Kong   2 Renmin University of China 3 Shenzhen University   4 City University of Hong Kong 1zzma2@cse.cuhk.edu.hk2wangwenxuan@ruc.edu.cn3wentichen7-c@my.cityu.edu.hk
1 Introduction
Large language models (LLMs) are demonstrating significant potential in healthcare, leading to a proliferation of benchmarks designed to evaluate their capabilities (Jin et al., 2021; Pal et al., 2022). These evaluation tools have evolved from early exam-style question answering (QA) to encompass more complex clinical tasks like report summarization and diagnosis (Wu et al., 2025b; Liu et al., 2024a; Wu et al., 2025a).
Figure 1: (a) Overall and (b) phase-by-phase performance of medical LLM benchmarks. The significant variance, identified through our MedCheck evaluation, highlights the core motivation for this work: the development of a principled framework for systematic benchmark evaluation.
However, despite their widespread adoption, the reliability of these benchmarks is a growing concern. Many evaluations rely heavily on closed-form, multiple-choice questions (MCQA) that, while testing factual knowledge, fail to assess open-ended clinical reasoning or account for real-world variability (Wu et al., 2025a; Zhang et al., 2025). Furthermore, a significant number of benchmarks are constructed from academic materials rather than authentic clinical data, which raises critical concerns about their clinical fidelity and the risk of data contamination—where models are evaluated on data seen during training (Ouyang et al., 2024; Wu et al., 2025b). These systemic flaws can create an illusion of progress, where models are optimized for tasks that lack genuine clinical utility.
Recent efforts have sought to improve evaluation quality by proposing general frameworks for AI benchmarks, such as BetterBench for general-purpose AI (Reuel et al., 2024) and How2Bench for code (Cao et al., 2025). While these frameworks have uncovered systemic issues in data quality and reproducibility, they are not tailored to the unique, high-stakes requirements of medicine. The medical domain demands a more rigorous, context-aware approach that accounts for specialized terminology, the ethical handling of sensitive patient data, and the paramount importance of safety and reliability.
To develop trustworthy clinical AI, the field must therefore transition from an ad-hoc, publication-driven approach to a disciplined, engineering-oriented paradigm for evaluation (Laskar et al., 2024; Yan et al., 2024). Mature fields like software engineering have long relied on structured methodologies such as the Software Development Life Cycle (SDLC) to manage complexity and ensure quality (Yas et al., 2023; Royce, 1970). Adopting a similar lifecycle-aware perspective is not merely beneficial for the science of AI evaluation; it is an absolute prerequisite for ensuring safety and efficacy in the medical domain (Park et al., 2020).
In this paper, we introduce MedCheck, the first comprehensive, lifecycle-oriented assessment framework specifically engineered for medical LLM benchmarks. MedCheck deconstructs benchmark development into five continuous stages, from design to governance, and provides a checklist of 46 medically-tailored criteria. To demonstrate its utility and diagnose the state of the field, we applied MedCheck in an in-depth empirical evaluation of 53 existing medical benchmarks. Our analysis reveals widespread, systemic issues: a profound disconnect from clinical practice, a crisis of data integrity from unmitigated contamination risks, and a systematic neglect of safety-critical dimensions like model robustness and uncertainty awareness.
Our contributions are threefold:
  • We introduce MedCheck, the first comprehensive, lifecycle-oriented evaluation framework with 46 criteria specifically designed for medical benchmarks.
  • We conduct an in-depth evaluation of 53 medical benchmarks, revealing widespread, systemic weaknesses across the current evaluation landscape.
  • We offer MedCheck as a practical checklist to guide the development of more reliable, transparent, and clinically relevant benchmarks for AI in healthcare.
2 Related works
2.1 The Evolution of Medical LLM Benchmarks
Evaluation of medical LLMs has been dominated by benchmarks derived from medical qualification examinations (Jin et al., 2021; Pal et al., 2022; Liu et al., 2023) and scholarly literature (Jin et al., 2019; He et al., 2020). While useful for assessing foundational medical knowledge, their format and content often lack direct clinical applicability.
To bridge this gap, recent efforts have focused on creating benchmarks that better mirror real-world clinical tasks. These include comprehensive benchmarks covering report summarization, diagnosis, and treatment planning (Wu et al., 2025b; Liu et al., 2024a; Wu et al., 2025a), as well as agentic benchmarks that simulate sequential clinical decision-making (Jiang et al., 2025; Schmidgall et al., 2024). Although these newer benchmarks represent a significant step forward, their development has often outpaced the creation of rigorous standards for their own evaluation, a gap our work aims to fill.
2.2 Towards a Science of Benchmark Evaluation
The proliferation of benchmarks has ignited a critical discussion about their reliability and validity. Alaa et al. (2025) empirically demonstrated a poor correlation between LLM performance on existing benchmarks and in real-world clinical scenarios, highlighting a crisis in the construct validity of current evaluation methods—the degree to which a test measures what it claims to measure (Cronbach and Meehl, 1955). This has catalyzed a new research area focused on the science of benchmarking itself.
Figure 2: The proposed five-phase lifecycle for medical benchmark engineering. This model illustrates the interconnected and continuous stages of development, from initial design and conceptualization to long-term documentation and governance.
Pioneering works have proposed general frameworks to standardize the evaluation of AI benchmarks. For instance, Bhardwaj et al. (2024) developed a checklist for data curation best practices. More comprehensively, BetterBench (Reuel et al., 2024) provides a 46-criteria framework for general-purpose AI benchmarks, and How2Bench (Cao et al., 2025) offers a 55-item checklist for code-related LLM benchmarks. These frameworks promote a lifecycle-aware perspective, revealing common issues in data quality, reproducibility, and transparency.
However, these valuable solutions are not specialized for medicine. High-stakes clinical applications demand a more rigorous, context-aware approach that accounts for specialized terminology, patient data ethics, and the paramount importance of safety and reliability. While existing surveys have profiled the medical LLM landscape (Liu et al., 2024c; Yan et al., 2024), they stop short of providing a structured framework for assessing benchmark quality. Our work bridges this critical gap by introducing MedCheck, the first assessment framework specifically engineered for the unique complexities of the medical domain.
3 Design
To systematically deconstruct the complexities and shortcomings of existing medical LLM benchmarks, we establish a conceptual framework for their development. This section introduces a novel five-phase lifecycle model for benchmark engineering and details the rigorous methodology employed in our comprehensive analysis.
3.1 The Five-Phase Medical Benchmark Lifecycle
Our proposed lifecycle, as depicted in Figure 2, provides a structured paradigm for the engineering of high-quality medical benchmarks. Each phase addresses a core set of objectives and potential pitfalls.
Phase I: Design and Conceptualization.
This foundational phase moves beyond mere task definition to establish the benchmark’s theoretical underpinnings, focusing on its construct validity—the degree to which it accurately measures its intended theoretical construct, such as "clinical reasoning" (Cronbach and Meehl, 1955; Mehandru et al., 2025; Schmidgall et al., 2024). As forcefully argued by Alaa et al. (2025), high performance on exam questions may not correlate with real-world clinical utility. This phase, therefore, defines the benchmark’s specific, clinically relevant purpose and its intended contribution, setting the standard against which its validity will later be judged.
Figure 3: The three-step methodology employed in this study. Our approach involved (1) developing the 46-criteria assessment framework, (2) systematically curating and scoring 53 benchmarks against it, and (3) performing a quantitative synthesis to generate systemic insights.
Phase II: Dataset Construction and Management.
This phase constitutes the benchmark’s empirical core, focusing on the curation of authentic, diverse, representative, and ethically sourced data. This demands strict adherence to medical privacy regulations like HIPAA and GDPR (Bhardwaj et al., 2024; Eke and Shuib, 2025). A primary challenge is proactively addressing data contamination, where evaluation data has been seen by an LLM during training (Ouyang et al., 2024; Wu et al., 2025b). Such contamination leads to inflated performance scores and misleading leaderboards (Dong et al., 2024; Cheng et al., 2025), making rigorous processes for data quality assurance, de-identification, and contamination detection paramount.
Phase III: Technical Implementation and Evaluation Methodology. This operational phase transforms a dataset into an evaluation toolkit with reproducible scripts and, crucially, metrics beyond simple accuracy (Chang et al., 2023). As closed-form MCQA may not assess deep reasoning (Wu et al., 2025a; Molfese et al., 2025), it is vital to evaluate the logical coherence of a model’s reasoning (Dai et al., 2025), its robustness to noisy inputs (Han et al., 2024), and its capacity to articulate uncertainty—a cornerstone of safe clinical practice (Shelmanov et al., 2025).
Phase IV: Benchmark Validity and Performance Verification. This phase provides the empirical validation for the benchmark as a measurement instrument. It involves presenting evidence for content validity (the content is representative of the clinical domain) and, critically, substantiating the construct validity claims established in Phase I (Alaa et al., 2025). A validated benchmark must also demonstrate its ability to reliably differentiate between models of varying capabilities, thus providing a clear signal of progress in the field (Reuel et al., 2024).
Phase V: Documentation, Openness, and Governance. This community-facing phase ensures the benchmark’s long-term value and impact. It involves clear documentation and open-source principles for transparency and reproducibility (Arvan et al., 2022; Cohen et al., 2018). Critically, it also demands a robust governance model for long-term maintenance, versioning, and community feedback to prevent the benchmark from becoming a static, decaying artifact.
3.2 Study Design
Our study employed a meticulous, multi-step methodology, illustrated in Figure 3, to ensure objectivity, reproducibility, and depth.
Step 1: Framework Development. We developed the 46-criteria MedCheck framework, which serves as the analytical lens for this study. The creation of this framework was a multi-stage process. First, we conducted an extensive literature review of existing best practices in general AI benchmarking (Reuel et al., 2024; Cao et al., 2025) and machine learning data curation (Bhardwaj et al., 2024). Second, we distilled these principles into a draft checklist. Finally, this checklist was iteratively refined into the final 46-criteria framework, grounded in the unique ethical and practical constraints of the medical domain, such as patient privacy (e.g., HIPAA compliance), the need for evidence-based standards, and the high stakes associated with potential patient harm (Farhud and Zokaei, 2021; Eke and Shuib, 2025). The complete 46-criteria checklist is provided in Appendix C.
Step 2: Systematic Benchmark Curation and Analysis. We curated and analyzed a corpus of 53 medical LLM benchmarks, listed in Appendix A. The selection process involved a systematic search for papers published between 2018-2025 using keywords such as "medical LLM," "medical benchmark," and "healthcare," aiming for comprehensive coverage of tasks, languages, and publication years. To ensure objectivity, a rigorous scoring protocol was implemented. Each benchmark was independently scored by a team of three NLP researchers with clinical informatics experience against the 46 criteria on a 3-point Likert scale (0: not met; 1: partially met; 2: fully met). All assessments were based exclusively on publicly available artifacts, including published papers, code repositories, and official websites. Any scoring discrepancies were resolved through a consensus discussion to arrive at a final score for each criterion.
Step 3: Quantitative Synthesis and Insight Generation. The individual scores were aggregated to enable a multi-level quantitative analysis. We calculated per-criterion average scores to identify specific, widespread weaknesses across the field. These were then rolled up into per-phase scores to assess the maturity of each stage in the development lifecycle. Finally, an overall score was computed for each benchmark to gauge its overall quality. This quantitative synthesis allowed us to move beyond anecdotal critiques and identify systemic, statistically significant deficiencies across the entire landscape of medical LLM evaluation, forming the evidence base for the guidelines presented in the following section. A detailed breakdown of these quantitative results can be found in Appendix B.
4 MedCheck: A Guideline for Medical Benchmarks
Based on our five-phase lifecycle and extensive analysis, we present the MedCheck framework as an actionable guideline. For each phase, we explain its core principle and summarize our key findings, further detailed in Appendix B.
Table 1: MedCheck Criteria for Phase I: Design and Conceptualization
No. Description
1 Does the benchmark define the targeted LLM capabilities in medicine for evaluation (e.g., QA, diagnostic reasoning)?
2 Does it describe specific clinical or research applications and their potential value?
3 Does it highlight its unique contribution or innovation compared to existing benchmarks?
4 Does it define the specific LLM functions for evaluation?
5 Does it define the scope of the medical specialties of the benchmark?
6 Is it designed to meet the needs of LLM researchers and clinical challenges?
7 Have qualified medical experts been involved in benchmark development?
8 Does it rely on recognized medical sources (e.g., clinical guidelines, databases)?
9 Does it align with international medical standards (e.g., ICD, SNOMED CT, LOINC)?
10 Are the metrics clear and closely related to clinical tasks?
11 Does it assess multiple dimensions (e.g., safety, completeness, interpretability)?
12 Does it consider potential risks and biases in model outputs?
4.1 Guideline for Design & Conceptualization
Explanation. This foundational phase anchors a benchmark in clinical relevance and scientific novelty. It necessitates the precise definition of the targeted LLM capabilities, medical scope, and application value, all of which must be grounded in documented domain expertise and authoritative sources. A robust design requires alignment with established medical standards (e.g., ICD), the use of clearly defined, multi-dimensional metrics that extend beyond accuracy, and the proactive consideration of safety and bias from inception (Arora et al., 2025).
Findings: The Clinical Disconnect. Our analysis reveals that while nearly all benchmarks (98%) define high-level objectives, this superficial clarity masks a systemic issue we term the Clinical Disconnect. A concerning 53% (28 of 53) fail to align with any formal medical standards (e.g., ICD, SNOMED CT), isolating them from the structured language of clinical practice. Furthermore, 47% (25 of 53) do not incorporate safety and fairness into their design, and 36% (19 of 53) evaluate only a single dimension like accuracy, neglecting critical aspects such as completeness. This disconnect stems from an "academic-first, clinical-second" mindset, where developers favor convenient data sources like exam questions from MedQA (Jin et al., 2021) and MedMCQA (Pal et al., 2022) over data reflecting complex clinical workflows. This convenience-driven design compromises clinical fidelity and construct validity (Alaa et al., 2025), leading to models optimized for irrelevant tasks—a perilous path that produces tools unfit and potentially unsafe for real-world deployment.
Table 2: MedCheck Criteria for Phase II: Dataset Construction & Management
No. Description
13 Are the original sources clearly stated and traceable?
14 Are the sources authoritative and well-justified?
15 Is the data origin clear, and is synthetic data validated?
16 Is the dataset representative of the target population?
17 Are diversity goals defined with supporting quantitative analysis?
18 Is the data properly cleaned and standardized?
19 Are privacy measures described and regulation-compliant?
20 Is the data format consistent and unambiguous?
21 Is expert-involved data review in place?
22 Are reference answers accurate and validated?
23 Are contamination risks detected and handled?
4.2 Guideline for Dataset Construction & Managemen
Explanation. This phase addresses the integrity of the benchmark’s core asset: its data. This principle mandates the use of traceable and authoritative data sources, with validated authenticity for any synthetic data. The dataset must be demonstrably representative and diverse, supported by quantitative analysis. Rigorous quality control—encompassing cleaning, standardization, expert review, and validated reference answers—is imperative, as are regulation-compliant privacy measures and the proactive mitigation of data contamination risks (Ouyang et al., 2024; Wu et al., 2025b).
Findings: A Crisis of Foundational Validity. Our analysis reveals critical weaknesses in data management that undermine the field’s empirical foundation. While most benchmarks are reasonably transparent about their primary data sources, subsequent quality control is severely lacking. A staggering 92% (49 of 53) fail to address data contamination, often by directly reusing data from other benchmarks without verification. Furthermore, 70% (37 of 53) are insufficiently diverse or representative, and 58% (31 of 53) lack any clear data audit or expert review mechanism. This widespread failure to ensure data integrity poses a direct threat to the field’s validity. Unvetted data may contain factual errors, while contamination leads to artificially inflated scores, creating a false perception of model capability (Magar and Schwartz, 2022; Deng et al., 2024; Bender et al., 2021). This renders leaderboards misleading and undermines the scientific credibility of reported results (Holistic AI, 2024), risking the construction of a field built upon unreliable evidence.
Table 3: MedCheck Criteria for Phase III: Technical Implementation & Evaluation Methodology
No. Description
24 Is the evaluation tool easy to install and use?
25 Are detailed documentation and environment settings provided to support reproducibility?
26 Are baseline models or human performance results provided for comparison?
27 Are there evaluations for the model’s reasoning process?
28 Are there evaluations testing the model’s robustness (e.g., input perturbations)?
29 Does the benchmark design help evaluate the generalization ability of models to unseen data?
30 Are there evaluations testing the model’s ability to express uncertainty?
31 Does the benchmark support both closed-source APIs and open-source models?
4.3 Guideline for Technical Implementation & Evaluation Methodology
Explanation. This phase focuses on transforming a dataset into a complete evaluation toolkit. It requires the provision of accessible, reproducible, and well-documented code that supports diverse model types and offers performance baselines for context. Fundamentally, the methodology must transcend accuracy to assess safety-critical capabilities: the model’s reasoning process, its robustness to input variations, its generalization to unseen data, and its capacity to articulate uncertainty—all of which are vital for trustworthy medical AI (Dai et al., 2025; Qiu et al., 2025; Han et al., 2024; Shelmanov et al., 2025).
Findings: Systematic Neglect of Safety-Critical Capabilities. This phase is the most underdeveloped in our analysis (average score: 52.4%), revealing a profound gap between current practice and the requirements for reliable medical AI. An alarming 94% (50 of 53) of benchmarks have no mechanism to test for model robustness, and 96% (51 of 53) fail to evaluate a model’s ability to handle uncertainty. Furthermore, 51% (27 of 53) neglect the model’s reasoning process, focusing only on the final answer. These omissions constitute a systematic neglect of safety. A model’s reasoning, robustness, and uncertainty awareness are cornerstones of clinical trustworthiness (Farhud and Zokaei, 2021). A brittle model that fails with slight data variations is dangerous (The BMJ, 2024), and an overconfident model that gives incorrect advice without expressing uncertainty is a direct threat to patient safety. By not measuring these capabilities, the community implicitly deems them unimportant, increasing the risk of deploying brittle, opaque, and unsafe systems.
Table 4: MedCheck Criteria for Phase IV: Benchmark Validity & Performance Verification
No. Description
32 Does the benchmark cover the claimed medical knowledge and skills?
33 Do the tasks realistically simulate clinical settings?
34 Can it distinguish between models of different levels?
35 Are benchmark scores correlated with real-world clinical performance?
36 Does the benchmark demonstrate internal consistency to ensure that different components reliably assess the same capability?
37 Are statistical tests used to verify results?
4.4 Guideline for Benchmark Validity & Performance Verification
Explanation. This phase concerns the scientific validation of the benchmark as a measurement instrument. It requires empirical evidence supporting both content validity (comprehensive coverage of the claimed domain) and construct validity (realistic task simulation and accurate measurement of the intended capability (Cronbach and Meehl, 1955)). A validated benchmark must also demonstrate its utility through proven discriminative power, correlation with real-world clinical performance, high internal consistency, and the application of statistical tests to verify results (Reuel et al., 2024).
Findings: The Risk of Misdirected Progress. Formal validation of a benchmark’s own scientific properties is rare. While most benchmarks can differentiate models, only 57% (30 of 53) provide a compelling analysis of their content validity, and just 40% (21 of 53) are grounded in scenarios with high real-world authenticity. Without proper validation, a benchmark’s results are difficult to interpret. Poor content validity creates blind spots, while poor construct validity means the benchmark may not measure the claimed capability at all, leading to misleading conclusions and misdirected research efforts (Alaa et al., 2025). This lack of self-reflection compromises the scientific value of the work built upon these tools, risking optimization for metrics that lack genuine clinical utility.
Table 5: MedCheck Criteria for Phase V: Documentation, Openness, and Governance
No. Description
38 Is there clear, comprehensive benchmark documentation?
39 Are the evaluation criteria and instructions clear and easy to follow?
40 Are limitations and potential risks openly discussed?
41 Has the benchmark undergone formal academic peer review?
42 Are the code and data publicly available with proper licensing?
43 Is a clear usage and citation guideline provided?
44 Is there a clear plan for updates and version control?
45 Is there an public channel for user feedback?
46 Is the long-term maintenance responsibility clearly stated?
4.5 Guideline for Documentation, Openness, & Governance
Explanation. This final phase ensures a benchmark’s long-term utility and trustworthiness. It requires comprehensive and transparent documentation, including clear instructions and a candid discussion of limitations. Adherence to open-source principles—publicly accessible code and data with appropriate licensing, clear citation guidelines, and academic peer review—is crucial. A robust governance model, which specifies long-term maintenance responsibilities, versioning plans, and a public feedback channel, is essential for sustained relevance and impact.
Findings: A Fragmented and Unsustainable Ecosystem. While most benchmarks provide public access to their assets (52 of 53), the governance required for sustainable impact is critically lacking. A significant 42% (22 of 53) do not specify a usage license, creating barriers to adoption. Most critically, 85% (45 of 53) have no clear maintenance plan, and 66% (35 of 53) lack a public feedback channel. This effectively renders them "fire-and-forget" artifacts destined for obsolescence. This practice fosters a fragmented and unsustainable ecosystem, where effort is wasted on disposable artifacts rather than on building a lasting, reliable infrastructure for scientific progress (Arvan et al., 2022).
5 Discussion
Our analysis of 53 medical LLM benchmarks, viewed through the MedCheck framework, reveals a field at a critical crossroads. While research is active, systemic deficiencies threaten the validity and utility of this work.
5.1 Implications: The Urgent Need for a Paradigm Shift
Continuing with the status quo is not only scientifically unsound but also clinically irresponsible. The current trajectory of medical LLM benchmarking fosters an "illusion of progress," where leaderboard scores climb on contaminated or clinically irrelevant tasks, masking a lack of genuine advancement in capabilities required for real-world application (Laskar et al., 2024; Chang et al., 2023). This misdirects research efforts, misinforms stakeholders, and delays the responsible integration of AI into healthcare by creating brittle, biased systems that could harm patients (Obermeyer et al., 2019).
A paradigm shift is urgently needed—a move from ad-hoc dataset creation toward a disciplined, lifecycle-aware practice of benchmark engineering. This requires treating benchmarks not as disposable artifacts for a single paper, but as scientific instruments demanding rigorous design, validation, and maintenance (Reuel et al., 2024).
5.2 Our Guideline as a Catalyst for Change
The five-phase lifecycle and 46-criteria checklist presented in this paper provide a practical toolkit for this essential shift. Our framework offers a clear, actionable roadmap for developers to create higher-quality benchmarks and for users to critically assess the limitations of existing ones. Adopting this framework can directly mitigate the critical issues identified in our study by promoting:
Clinical Grounding: Mandating expert involvement and alignment with medical standards from Phase I.
Data Integrity: Enforcing transparent sourcing, quality control, and contamination checks in Phase II.
Safety-Oriented Evaluation: Requiring the assessment of reasoning, robustness, and uncertainty in Phase III.
Scientific Validity: Emphasizing content and construct validation in Phase IV.
Sustainable Impact: Promoting open practices and long-term governance in Phase V.
5.3 Future Directions: The Next Frontier of Medical Benchmark Research
Our work illuminates a clear path forward. Future efforts must move beyond the static, multiple-choice paradigm that currently dominates and pursue three critical directions:
Embracing Dynamic and Interactive Benchmarks. Future benchmarks must reflect the dynamic nature of clinical encounters by assessing sequential decision-making and information gathering. Pioneering works like MediQ (Li et al., 2024), which evaluates question-asking, and AgentClinic (Schmidgall et al., 2024), which uses multi-agent simulation, exemplify this necessary evolution toward evaluating core clinical reasoning skills (Trimble and Hamilton, 2016).
Prioritizing Empirical Construct Validity. The field must empirically validate that benchmarks truly measure the clinical constructs they claim to (Alaa et al., 2025). This requires innovative methodologies that correlate benchmark scores with performance on real-world clinical data, such as predicting patient outcomes from EHRs. Without this validation, we risk optimizing for metrics that lack clinical significance.
Building a Collaborative Evaluation Ecosystem. To foster quality and transparency, the community should develop a living repository, akin to platforms like betterbench.stanford.edu (Reuel et al., 2024). Such a platform would enable continuous evaluation of benchmarks against a standardized framework like MedCheck. This would create a positive feedback loop, incentivizing higher-quality benchmarks and empowering all stakeholders to make more informed decisions about developing and deploying medical LLMs.
6 Conclusion
To address the lack of reliability in medical benchmarks, we introduced MedCheck, a lifecycle-oriented framework with 46 criteria. Our evaluation of 53 benchmarks revealed systemic deficiencies, including a "Clinical Disconnect" from real-world practice, a "Crisis of Data Integrity" due to contamination risks, and a "Systematic Neglect of Safety-Critical Capabilities." These issues underscore the need to shift from ad-hoc dataset creation to a more disciplined approach. MedCheck provides the essential toolkit for this shift, guiding the development of a rigorous infrastructure to move beyond the leaderboard and rethink medical LLM benchmarks, enabling the creation of genuinely safe and effective clinical AI.
Limitations
We acknowledge this study’s limitations. First, our analysis of 53 benchmarks, while extensive, is not exhaustive given the field’s rapid growth. Second, the scoring process, despite a rigorous protocol, retains a degree of subjectivity inherent in qualitative assessment. Third, our findings are based exclusively on public artifacts, which may not capture all unpublished development practices. Finally, the MedCheck framework itself is a snapshot of current best practices and will require future revisions to keep pace with evolving AI capabilities and ethical standards.
Acknowledgments
Appendix
In this appendix, we provide additional information about MedCheck. Appendix A lists the 53 evaluated medical LLM benchmarks. Appendix B provides extra quantitative analysis of the evaluation results. Appendix C shows the complete list of evaluation criteria.
Appendix A List of Evaluated Benchmarks
We assessed the following 53 medical LLM benchmarks (alphabetical order):
  • AfriMed-QA Olatunji et al. (2025)
  • AgentClinic Schmidgall et al. (2024)
  • Asclepius Liu et al. (2025b)
  • BRIDGE Wu et al. (2025b)
  • CHBench Guo et al. (2024)
  • CheXpert Irvin et al. (2019)
  • CLIMB Dai et al. (2025)
  • CliMedBench Ouyang et al. (2024)
  • ClinicBench Liu et al. (2024a)
  • CMB Wang et al. (2024a)
  • cMedQA2 Zhang et al. (2018)
  • CMExam Liu et al. (2023)
  • COGNET-MD Panagoulias et al. (2024)
  • DataDEL Yang et al. (2024)
  • EHRNoteQA Kweon et al. (2024)
  • EndoBench Liu et al. (2025c)
  • ER-REASON Mehandru et al. (2025)
  • GMAI-MMBench Chen et al. (2024)
  • HeadQA Vilares and Gómez-Rodríguez (2019)
  • HealthBench Arora et al. (2025)
  • LLMEval-Med Zhang et al. (2025)
  • MedAgentBench Jiang et al. (2025)
  • MedAgentsBench Tang et al. (2025)
  • MedCalc Khandekar et al. (2024)
  • MedChain Liu et al. (2024b)
  • MedConceptsQA Shoham and Rappoport (2024)
  • MEDEC Abacha et al. (2025)
  • MedExQA Kim et al. (2024)
  • MEDIC Kanithi et al. (2024)
  • MediQ Li et al. (2024)
  • MedJourney Wu et al. (2025c)
  • MedMCQA Pal et al. (2022)
  • MedOdyssey Fan et al. (2025a)
  • MedQA Jin et al. (2021)
  • MedR-Bench Qiu et al. (2025)
  • MedRisk Liu et al. (2025a)
  • MedSafetyBench Han et al. (2024)
  • MedS-Bench Wu et al. (2025a)
  • MLEC-QA Li et al. (2021)
  • MMMU (Health & Medicine) Yue et al. (2024)
  • MVME (AI hospital) Fan et al. (2025b)
  • OmniMedVQA Hu et al. (2024)
  • PathMMU Sun et al. (2025b)
  • PathVQA He et al. (2020)
  • PMC-VQA Zhang et al. (2024)
  • PubMedQA Jin et al. (2019)
  • ReasonMed Sun et al. (2025a)
  • SLAKE Liu et al. (2021)
  • TrialPanorama Wang et al. (2025)
  • VQA-Med Ben Abacha et al. (2021)
  • VQA-RAD Lau et al. (2018)
  • webMedQA He et al. (2019)
  • XMedBench Wang et al. (2024b)
Appendix B Quantitative Results
This appendix provides a summary of the quantitative analysis conducted on the 53 medical LLM benchmarks.
B.1 Publication Year and Task Focus of Benchmarks
Our collection covers open-source medical large model benchmarks published between 2018 and 2025.
Annual Trend. As shown in Figure 4, the publication years of the 53 benchmarks we collected. A clear turning point can be observed in 2024, when the volume of medical benchmark publications significantly increased. Prior to 2024, only 15 benchmarks were released. The brief surge in 2019, driven by the rise of deep learning and Transformer models, catalyzed short-term growth in medical AI and a peak in benchmark publications. However, no substantial upward trend was evident before 2024.
Figure 4: Publication Year Distribution of the 53 Medical Benchmarks
In contrast, with the emergence of large models such as ChatGPT and Med-PaLM in 2024, there was an explosive growth in medical NLP and multimodal benchmarks. A total of 26 medical benchmarks were published in 2024, surpassing the cumulative number of the previous six years. As of July 10, 2025, 12 new benchmarks were already released in the first half of 2025, showing a steady growth trend.
Task Trend. Before 2024, most benchmarks primarily focused on evaluating models’ capabilities in NLP and VQA. As shown in Figure 5, following the explosion in benchmark quantity post-2024, the range of tasks covered expanded to six categories, with new benchmarks addressing specific tasks for large models, such as multimodal understanding, reasoning and decision-making, dialogue/interactive QA , as well as medical text interpretation and safety in medical recommendation generation.
Figure 5: Temporal Distribution of Benchmarks by Task Type
In summary, as illustrated in Figure 6 nearly half (47%) of the existing medical benchmarks focus on assessing models’ text comprehension, NLP, and VQA abilities. This indicates that most medical benchmarks historically centered around structured or unstructured medical text understanding. However, in recent years, there has been an increasing trend towards medical question answering and clinical NLP, with the highest number of such benchmarks being published in the first half of 2025.
Figure 6: Task Type Distribution of the 53 Medical Benchmarks
B.2 Natural Language
Figure 7 presents the distribution of natural language usage across the 53 collected benchmarks. It is evident that English dominates, being used in 42 of the benchmarks (approximately 79%), establishing itself as the primary language for evaluating medical LLMs. Chinese ranks second, appearing in 17 benchmarks (approximately 32%), indicating a notable level of development in Chinese medical AI. In contrast, other languages such as Arabic, Russian, Portuguese, Japanese, and Hindi—despite being spoken by large global populations—remain underrepresented in publicly available benchmarks, with each language appearing in only 2 to 6 benchmarks on average. This highlights that multilingual medical AI remains in its early stages.
Figure 7: Natural Language Distribution of the Benchmarks
Finally, it may be concluded that current medical LLM benchmarks exhibit a pronounced trend of language centralization. English, as the de facto language of international medical communication, dominates the vast majority of evaluation tasks. Although multilingual benchmarks are still in their infancy, they hold substantial potential in promoting global medical equity and cross-cultural adaptation. Moving forward, greater efforts are needed to improve the availability of multilingual data resources and to develop cross-lingual evaluation strategies.
B.3 Medical Diversity
Distribution of Diseases. Figure 8 illustrates the distribution of disease categories represented in the 53 medical benchmarks, categorized according to the first 23 chapters of the ICD-11 classification system (with Certain conditions originating in the perinatal period and Developmental anomalies excluded due to the absence of relevant benchmarks). The analysis reveals that current research on medical LLMs is heavily concentrated in disease areas such as Diseases of the musculoskeletal system or connective tissue (16 benchmarks), Diseases of the skin (15), Diseases of the respiratory system (14), and Neoplasms (13). These categories typically have abundant publicly available imaging data (e.g., X-rays, CT scans, pathology slides), which facilitates the development of standardized evaluation protocols based on both image and text modalities.
Figure 8: Disease Distribution Based on ICD-11 Classification
In contrast, low-frequency disease categories such as Mental and behavioral disorders (e.g., depression, anxiety) rely on psychological scales, clinical interviews, or life history data, which are difficult to standardize and rarely captured in benchmark datasets. Similarly, Immunological and Hematological diseases often require the integration of laboratory indicators, clinical progression, and imaging, making them difficult to model using unimodal approaches. From a clinical perspective, however, these underrepresented diseases are not necessarily rare in real-world settings. Many exhibit high outpatient visit rates, cause significant patient distress, or carry substantial social burden. This highlights a misalignment between research frequency and clinical demand. Future research in medical LLMs should be designed with greater awareness of this gap, ensuring that high-demand but rarely covered disease populations are not systematically neglected.
Medical Specialties. While ICD-11 provides a useful framework for disease-based diversity analysis, it lacks granularity in distinguishing among interdisciplinary or population-specific fields such as clinical medicine, obstetrics and gynecology, and pediatrics. To complement ICD-11-based analysis, we additionally introduce a second diversity metric based on the official list of clinical departments issued by the National Health Commission of China. Figure 9 shows the distribution of benchmarks across these medical specialties (excluding departments such as Endemic Diseases that are not represented in any benchmark or that overlap with ICD-11 categories).
Figure 9: Medical Specialty Distribution Based on the Official Directory of Clinical Departments
This classification reveals a similar pattern: departments with well-structured data and clearly defined tasks dominate the landscape, such as Surgery (12 benchmarks), General Practice (11), and Medical Imaging (9). In contrast, specialties associated with vulnerable populations or complex care—such as Rehabilitation Medicine, Occupational Health, and Child Health Care—are significantly underrepresented.
This imbalance may result in limited generalizability and fairness in real-world deployment of medical LLMs, particularly in addressing the needs of underserved or structurally marginalized groups. Therefore, future benchmark design should aim for more balanced coverage across medical specialties and actively promote the standardization and open sharing of data from underrepresented departments. This will support the development of AI systems that are not only technically robust but also clinically inclusive.
B.4 Benchmark Performance Heatmap Example
Figure 10: Performance score rates (%) for 14 medical LLM benchmarks across the five phases of the MedCheck evaluation framework.
The heatmap in Figure 10 visualizes the performance of 14 out of 53 of the medical LLM benchmarks evaluated evaluated against the five-phase MedCheck framework. Each row represents a benchmark, and columns correspond to the framework’s phases. The color intensity reflects the score rate (%), highlighting the comparative strengths and weaknesses of each benchmark and revealing systemic trends.
B.5 Statistics About Phase I: Design and Conceptualization
This section presents a systematic evaluation of the design logic and conceptual clarity of existing medical benchmarks. we examine their alignment with medical capabilities, practical application scenarios, and innovation objectives. The assessment is structured around three core aspects: (1) the clarity of domain positioning from the perspective of AI capabilities and clinical scope, (2) the methodological soundness in terms of metric design and evaluation diversity, and (3) the integration of safety and fairness considerations as part of responsible AI practices.
As shown in Figure 11, the average normalized score across benchmarks is 74.3% (17.8 out of 24), with a mean of 1.5 points per criterion, indicating general compliance with baseline expectations. However, performance on several key items remains suboptimal and warrants further attention.
Figure 11: Scoring Performance of the 53 Benchmarks in the Design and Conceptualization Stage
Clearly specified application scenarios. To ensure practical relevance, benchmarks should clearly define their intended clinical or biomedical research scenarios. Such grounding facilitates real-world applicability and aligns evaluation with deployment needs. As illustrated in Figure 12, 66% of benchmarks specify concrete application scenarios, articulating clinical value and expected utility. In contrast, 19% provide only general references to decision-support contexts, and 15% fail to specify any scenario, reducing interpretability and downstream usability.
Figure 12: Performance on Clarity of Application Scenario
Involvement of domain experts. Expert input is a critical component in medical AI development. It ensures the appropriateness of task design, data annotation quality, and answer validity. However, as shown in Figure 13, only 25% of benchmarks explicitly report domain expert involvement, often without detailing the roles or qualifications. 17% do not mention expert participation, raising concerns about the benchmarks’ clinical validity. Future benchmark design should adopt standardized expert reporting practices to enhance credibility, transparency, and methodological rigor.
Figure 13: Performance on Domain Experts Involvement
Alignment with medical standards. Adherence to international medical standards (e.g., ICD-11, SNOMED CT, LOINC) is essential for achieving semantic interoperability, annotation consistency, and cross-system comparability. As Figure 14 indicates, 53% of benchmarks do not clarify whether their data conforms to such standards. This omission undermines the benchmark’s reusability and weakens the comparability of evaluation outcomes. Without standard-based structure and terminology, label ambiguity and inconsistent categorization may arise, which compromises the reliability of model performance evaluation.
Figure 14: Performance on Medical Standards Alignment
Multidimensional evaluation design. In medical AI, a model’s performance cannot be solely measured by accuracy. Outputs that are correct but unsafe, incomplete, or biased may have detrimental consequences. Multidimensional evaluation frameworks are required to assess whether models’ output is not only correct but also reliable. As presented in Figure 15, 36% of benchmarks use only a single metric, typically accuracy, and 21% mention other metrics without detailed implementation. Metrics such as uncertainty estimation, answer refusal, and harmful output detection are absent from most benchmarks, despite their relevance in high-stakes medical decision-making.
Figure 15: Performance on Multi-dimensional Evaluation
Consideration of safety and fairness. Robust evaluation frameworks must account for safety and fairness, especially in sensitive domains like healthcare. As shown in Figure 16, only 15% of benchmarks include explicit safety or fairness evaluations, while 38% acknowledge their importance without implementation. The remaining 47% omit these aspects entirely. Neglecting safety assessments increases the risk of misleading or overconfident outputs, which may harm clinical decisions. The lack of fairness evaluation may propagate systematic bias across gender, age, race, or disease distribution, exacerbating healthcare disparities. To build trustworthy and ethical medical AI, safety and fairness must be treated as core evaluation dimensions.
Figure 16: Performance on Consideration of Safety and Fairness
B.6 Statistics About Phase II: Dataset Construction and Management
This phase constitutes the empirical foundation of the benchmark, with an emphasis on constructing datasets that are realistic, diverse, representative, and ethically sourced. Key elements include stringent procedures for data quality assurance, privacy protection, and contamination prevention. The dataset sources and quality were assessed across five dimensions: transparency and traceability of data provenance, reliability of data sources, and data authenticity. Data processing and privacy protection were evaluated in six aspects, including data cleaning and standardization, privacy-preserving mechanisms, and data formatting protocols. As shown in Figure 17, the 53 benchmarks achieved an average score rate of 74.6% (16.4 out of 22) across 11 evaluation criteria, with a mean of 1.4 points per item. Overall, the results indicate general compliance with the proposed standards, although two specific criteria showed suboptimal performance.
Figure 17: Scoring Performance of the 53 Benchmarks in the Phase II: Dataset Construction and Management
Dataset Diversity. This criterion evaluates whether the benchmark explicitly defines diversity objectives—such as coverage of disease types and clinical departments—and provides quantitative evidence (e.g., disease distribution histograms, task type ratios) to demonstrate dataset coverage. Ensuring sufficient diversity is critical to evaluate model generalization across a broad range of cases and specialties, thereby mitigating evaluation bias.
The specific calculation is shown in Equation  1. In this Equation, $N_{\text{disease}}$ represents the number of diseases in the ICD-11 standard, specifically the first 23 diseases, serving as the benchmark for the types of diseases considered. $N_{\text{department}}$ denotes the number of medical departments, typically referring to the medical specialties included in the model’s evaluation. The variable $N_{\text{disease}}^{\text{benchmark}}$ indicates the number of diseases covered by the benchmark, representing the actual disease categories included in the model’s evaluation. Similarly, $N_{\text{department}}^{\text{benchmark}}$ refers to the number of medical departments involved in the benchmark. Finally, $R_{\text{coverage}}$ represents the coverage ratio, which is calculated by comparing the sum of the diseases and medical departments covered by the benchmark $N_{\text{disease}}^{\text{benchmark}}+N_{\text{department}}^{\text{benchmark}}$ to the total number of diseases and medical departments in the reference standard $N_{\text{disease}}+N_{\text{department}}$. This formula provides a quantitative measure of the benchmark’s coverage across various diseases and medical departments, reflecting the model’s diversity and comprehensiveness in its evaluation.
The coverage of each benchmark after calculation is shown in Fig. 18. We scored each benchmark based on the average total coverage rate (21.8%), and the results are presented in Fig 19, 38% (20 out of 53) of the benchmarks lack a clear definition or discussion of dataset diversity. The absence of well-defined diversity goals and corresponding quantitative analyses hinders the ability to assess model robustness across different disease categories or clinical settings, potentially obscuring vulnerabilities in specific subgroups or task scenarios.
$R_{\text{coverage}}=\frac{N_{\text{disease}}^{\text{benchmark}}+N_{\text{department}}^{\text{benchmark}}}{N_{\text{disease}}+N_{\text{department}}}$ (1)
Figure 18: Coverage of disease types, medical departments, and overall performance across 53 medical benchmarks.
Figure 19: Performance on Dataset diversity
Prevention of Data Contamination. This dimension assesses whether the benchmark includes mechanisms to identify and mitigate potential data contamination—i.e., the presence of evaluation samples within the training corpus of LLMs. Such contamination can artificially inflate model performance through memorization, thereby compromising the fairness and validity of the evaluation. If present in high-risk tasks (e.g., diagnostic decisions or medication recommendations), contamination may mislead researchers and regulators in their assessment of model safety. Therefore, contamination detection is a prerequisite for ensuring the credibility and scientific rigor of benchmark evaluations. However, as depicted in Figure 20, only 8% (4 out of 53) of the benchmarks implemented both detection and mitigation strategies. Another 49% (26 out of 53) conducted detection without addressing identified contamination, while 43% (23 out of 53) did not mention the issue at all.
Figure 20: Performance on Data Contamination Prevention
B.7 Statistics About Phase III: Technical Implementation and Evaluation Methodology
This is the operational backbone. It consists of 8 items, encompassing the development of accessible, reproducible evaluation scripts and the selection of metrics that go beyond simple accuracy to assess deeper capabilities such as reasoning, robustness, and uncertainty awareness. For the 53 evaluated benchmarks, the average score is 52.4% (8.38 out if 16), with a mean score of 1.05 for each item. Results indicate that this is the second worst performing stage among the 5 stages. As shown in Figure 21, there are 3 items exhibiting critical underperformance, where majority of the benchmarks receive 0 mark.
Figure 21: Scoring Performance of the 53 Benchmarks in the Phase III: Technical Implementation and Evaluation Methodology
Reasoning Process Evaluation. This criterion assesses whether the benchmarks include evaluation for the reasoning process of the models. Apart from the final answer, understanding the decision-making process of models equally important.
As shown in Figure 22, a concerning 51% (27 out of 53) of benchmarks do not have any consideration for the models’ reasoning process. While 24% (13 out of 53) of benchmarks mention the importance of evaluating the reasoning process, only 25% (13 out of 53) of benchmarks design concrete assessment for reasoning process with clear methods and metrics.
Without evaluating the reasoning process, the flawed logic or hidden biases of models may left unchecked, potentially compromising patient safety. By assessing the reasoning path, how and why a model arrives at its conclusions can be revealed, fostering transparency and trustworthiness.
Figure 22: Performance on Reasoning Process Evaluation
Robustness Evaluation. This item assesses whether there is evaluation for robustness in the benchmark. In practical application, a robust model should be able to reliably interpret different variations of input. Robustness evaluation help assess whether models are resilient to imperfections and perturbations, ensuring the reliability when adopted in high-stakes medical scenarios.
As depicted in Figure 23, an overwhelming 87% (46 out of 53) of the covered benchmarks do not consider robustness in their evaluation. While 4 benchmarks recognize the importance of robustness evaluation, only 3 benchmarks design clear assessment for robustness.
Real-world environments are inherently variable and noisy. While models may perform well under idealized inputs in the benchmark, they may fail when facing subtle changes. Without robustness evaluation, benchmark results could be overly optimistic. This false confidence may lead to misleading decisions, causing risks when deployed in real-world environment. By evaluating robustness, users can gain a better understanding of models’ stability and consistency, fostering safer and more dependable deployment in complex medical environments.
Figure 23: Performance on Robustness evaluation
Uncertainty Evaluation. This criterion is designed for evaluating the ability of models to recognize and express its uncertainty. In high-stakes medical contexts, it is vital for models to recognize and communicate the limits of their knowledge. Assessing how safely and responsibly a model handles uncertainty promotes caution, reducing the risk of overconfident errors and misleading decisions.
As illustrated in Figure 24, 73% (39 out of 53) benchmarks have no considerations for uncertainty in their evaluation. Even though 23% (12 out of 53) acknowledge the importance of uncertainty evaluation, only 4% (2 out of 53) incorporate evaluation related to uncertainty in the benchmark, highlighting a significant gap in current benchmarking practices.
If uncertainty evaluation is omitted, it can lead to misleading performance results that overstate a model’s safety and reliability. When facing uncertainty, if models provide overconfident answers, it could lead to harmful outcomes. Benchmarks without uncertainty evaluation fail to reward models that behave ethically and responsibly under uncertainty, undermining the development of safer AI tools in medicine. By evaluating uncertainty, benchmarks can better reflect clinical uncertainty, fostering more trustworthy and safer decision-making behavior
Figure 24: Performance on Uncertainty Evaluation
B.8 Statistics About Phase IV: Benchmark Validity and Performance Verification
This phase provides scientific validation for the benchmark. It encompasses evidence of content validity (i.e., whether the benchmark comprehensively covers the target domain) and construct validity (i.e., whether it measures the claimed latent capabilities), as well as its ability to distinguish performance differences among models of varying proficiency levels. Among the 53 benchmarks evaluated, the average score rate for this phase was 48.8% (5.9 out of 12), with a mean of 0.5 points per item—representing the lowest performance among the five phases. As demonstrated in Figure 25, of the six evaluation criteria in this phase, three demonstrated notably poor results.
Figure 25: Scoring Performance of the 53 Benchmarks in the Phase IV: Benchmark Validity and Performance Verification
Correlation with Clinical Performance. This component assesses whether there is empirical evidence exploring the relationship between benchmark scores and model performance in real-world clinical applications. Specifically, it examines whether the benchmark scores can reliably reflect clinical effectiveness—an essential indicator of external validity and practical utility. In medical contexts, the ultimate goal of model performance is to enhance the accuracy, safety, and efficiency of clinical decision-making. If evaluation scores fail to map onto actual clinical outcomes, the benchmark may mislead researchers regarding model applicability. Moreover, neglecting this correlation may incentivize developers to over-optimize non-essential capabilities for better scores, thus deviating from real-world deployment needs.
As presented in Figure 26, among the 53 benchmarks reviewed, 45% (24 benchmarks) did not address the correlation between evaluation scores and clinical performance. A further 53% (28 benchmarks) discussed expected correlations at a theoretical level but lacked supporting experimental evidence. Only 2% (1 benchmark) provided preliminary empirical analysis exploring the correlation, along with a discussion of the results. These findings indicate a significant gap in validating the practical relevance of benchmark scores. Future medical benchmarks should prioritize establishing and verifying correlations between evaluation outcomes and clinical workflows or expert assessments. This is essential to ensure that benchmark scores serve as credible and informative indicators for real-world model deployment.
Figure 26: Performance on Correlation with Clinical Performance
Internal Consistency. This metric examines whether different components or items within a benchmark consistently measure the same target capability, thereby ensuring internal reliability. High internal consistency reduces the risk that variability in item quality, phrasing, or task context will distort evaluation results. Figure 27 shows that in this study, only 4% (2 out of 53) of the benchmarks reported statistical indicators of internal consistency, with positive results. An additional 24% (13 out of 53) also reported such metrics, but without clear discussion or yielded inconclusive findings. The remaining 72% (38 out of 53) did not conduct internal consistency assessments. The absence of such analysis raises concerns regarding evaluation accuracy and fairness. Inconsistent benchmarks may obscure true model strengths or weaknesses across specific skill dimensions, impeding targeted model refinement. In clinical applications, this may further compromise risk assessment and regulatory oversight. Therefore, future benchmarks should prioritize the evaluation and reporting of internal consistency to enhance reliability.
Figure 27: Performance on Internal consistency
Statistical Significance Reporting. This criterion evaluates whether statistical significance or uncertainty—such as confidence intervals or p-values—was reported when comparing model performances. In the absence of such analyses, observed performance differences may arise from sample randomness or dataset variability rather than genuine model capability disparities. This undermines the scientific rigor of the benchmark and may result in erroneous conclusions during model selection or optimization. In high-stakes medical scenarios, such inaccuracies pose safety risks if incorporated into decision-support systems or diagnostic tools. Consequently, the systematic reporting of statistical measures such as confidence intervals, p-values, or variance analyses is essential for constructing trustworthy and interpretable evaluation frameworks.
It can be seen from Figure 28 that, in practice, 59% (31 out of 53) of the benchmarks reported only single-run results. Although 32% (17 out of 53) provided mean and standard deviation across multiple runs, they did not employ formal statistical tests to compare model performance. These findings highlight a lack of rigorous statistical standards in current benchmark design. Incorporating structured multi-run strategies and statistical significance testing is critical to enhance the reliability and validity of performance evaluations in medical AI.
Figure 28: Performance on Statistical significance Reporting
B.9 Statistics About Phase V: Documentation, Openness, and Governance
Documentation, Openness, and Governance represent community-facing components that are essential for ensuring the long-term value and impact of a benchmark. This phase includes the provision of clear documentation, adherence to open-source principles, and the establishment of governance mechanisms for maintenance, version control, and community feedback. Among the 53 benchmarks evaluated, as depicted in Figure 29, the average score rate for this phase was 65.2% (11.7 out of 18), with a mean score of 1.0 per item. The scoring distribution in this phase was relatively balanced, with only one criterion showing slightly weaker performance.
Figure 29: Scoring Performance of the 53 Benchmarks in the Phase V: Documentation, Openness, and Governance
Discussion of Limitations and Risks. This criterion evaluates whether the benchmark documentation openly and explicitly discusses its inherent limitations, potential societal risks, or misuse scenarios. No benchmark is flawless. Proactively disclosing such limitations and risks reflects scientific rigor and a strong sense of responsibility on the part of the developers. It also helps prevent overreliance on benchmark results or inappropriate applications. Figure 30 shows that Only 7% (4 out of 53) of the benchmarks provided thorough and candid discussions of limitations and associated risks. A majority—74% (39 out of 53)—briefly mentioned limitations but lacked in-depth analysis. The remaining 19% (10 out of 53) did not discuss any form of limitation or risk.
In the medical domain, the responsibilities of a benchmark extend beyond merely providing performance scores. Benchmarks should also serve as tools for risk communication and usage guidance. Failure to include such information undermines both scientific robustness and standardization, and diminishes the benchmark’s value for clinical practice and policy-making. Therefore, systematic disclosure of limitations and potential risks should not be considered optional, but rather a fundamental requirement in the design and dissemination of medical benchmarks.
Figure 30: Performance on Discussion of Limitations and Risks
B.10 Scores per lifecycle Stage
We present the top 15 benchmarks for each of the five lifecycle phases using bar charts(Figure31 to Figure35). The scores for each phase reflect performance in areas such as design, dataset construction, technical implementation, validation, and documentation and governance. This visualization allows for a clear comparison of benchmark performance across phases, highlighting strengths in certain areas and identifying potential weaknesses in others.
Figure 31: Top 15 Benchmarks by Score in Phase I
Figure 32: Top 15 Benchmarks by Score in Phase II
Figure 33: Top 15 Benchmarks by Score in Phase III
Figure 34: Top 15 Benchmarks by Score in Phase IV
Figure 35: Top 15 Benchmarks by Score in Phase V
Appendix C Full Evaluation Criteria
C.1 Design and Conceptualization
C.1.1 Purpose and Intent
  1. 1.
    Clarity of Evaluation Objectives
    • Explanation: Benchmark developers should clearly specifies the targeted capabilities of LLMs it aims to evaluate in the medical field (e.g., medical knowledge question answering, diagnosis, report generation).
    • Justification: Clearly defined evaluation objectives can avoid ambiguity, facilitating subsequent data collection, task design, and metric selection. Also, it helps users determine whether the benchmark aligns with their specific evaluation needs.
    • Scoring:
      • 0: Does not mention or define the evaluation objective
      • 1: Mentions the evaluation objectives but the definition is broad or insufficiently detailed.
      • 2: Explicitly defines the medical capabilities being evaluated and provides examples.
  2. 2.
    Clarity of Application Scenario
    • Explanation: Benchmark developers should clearly describes the specific medical application scenarios it corresponds to and explains the potential value.
    • Justification: Linking the benchmark to real-world application scenarios ensures that the results are meaningful. It facilitates the validation of model’s effectiveness in specific settings, ultimately better serving users such as doctors and researchers.
    • Scoring:
      • 0: Does not describe any application scenarios.
      • 1: Mentions application scenarios but provides vague descriptions or fails to clarify the potential value.
      • 2: Clearly describes the application scenarios and elaborates on the value in detail.
  3. 3.
    Uniqueness and Novelty
    • Explanation: By comparing with relevant benchmarks, benchmark developers should explain their contributions and the uniqueness of the benchmark, such as filling a gap and proposing new evaluation methodology.
    • Justification: Demonstrating the uniqueness demonstrates the necessity and justification of the new benchmark. It also helps the community to better understand its unique value, promoting continual innovation in evaluation and benchmarking.
    • Scoring:
      • 0: Does not compare with relevant benchmarks or does not mention its uniqueness.
      • 1: Briefly mentions other benchmark, but the comparison is insufficient to explain its unique value.
      • 2: Provides detailed comparison with relevant benchmarks and clearly argues its uniqueness.
C.1.2 Scope and Applicability
  1. 1.
    Target Capability of Evaluation
    • Explanation: Benchmark developers should clearly define the capability of LLMs intended to evaluate (e.g., text generation, multimodal understanding)
    • Justification: Clearly defining the targeted capability helps clarify the scope of the benchmark, which can prevent misusing the benchmark.
    • Scoring:
      • 0: Does not define the target LLM capability.
      • 1: Briefly mentions the target capability without clear definition.
      • 2: Clearly defines and explains the target LLM capability.
  2. 2.
    Medical Domain Coverage
    • Explanation: Benchmark developers should clearly define the scope of the medical specialties of the benchmark, such as clinical departments, disease types, or task types.
    • Justification: By clearly defining the medical scope, it helps users better understand the breath and depth of the coverage of the benchmark. It also helps users determine whether the benchmark aligns with their needs.
    • Scoring:
      • 0: The medical scope and coverage are not defined.
      • 1: The medical scope and coverage are only briefly mentioned.
      • 2: The medical scope and coverage are clearly defined and explained.
  3. 3.
    Demonstration of User Needs
    • Explanation: The benchmark should reflect the core concerns and assessment needs of its target users, such as addressing specific clinical challenges or overcoming technical obstacles.
    • Justification: An effective benchmark should serve the needs of users. Evidence of user needs justifies the development of the benchmark. This boosts credibility, promotes adoption, and helps guide model improvements.
    • Scoring:
      • 0: Does not reflect any consideration of user needs.
      • 1: Briefly mentions user needs, but lacks concrete external evidence.
      • 2: Clearly explains user needs to justify design choices and connects user needs to benchmark tasks (e.g., referencing relevant literature, user surveys, or expert interviews).
C.1.3 Medical Expertise and Professionalism
  1. 1.
    Domain Experts Involvement
    • Explanation: Domain expert (e.g., physicians or clinical researchers) should be involved in the development of the benchmark.
    • Justification: Due to the professionalism and rigor required in the medical field, the development of a benchmark must involve deep engagement from domain experts. Their expertise are fundamental to ensuring data quality, task validity, authenticity and relevance.
    • Scoring:
      • 0: No mention of medical expert participation.
      • 1: Briefly mentions expert involvement, but the information is unclear and lacks details on the roles or involved tasks.
      • 2: Clearly describes the experts’ qualifications and their involvement, including specific roles and tasks throughout the development process.
  2. 2.
    Authoritative Knowledge Sources
    • Explanation: Benchmark developers must clearly specify which authoritative medical knowledge sources (e.g., clinical guidelines, textbooks, medical databases) the benchmark content is based on.
    • Justification: Benchmark content should adhere to recognized, evidence-based medical knowledge sources. By using authoritative sources, it ensures that the benchmark is scientifically reliable, enhancing credibility and transparency.
    • Scoring:
      • 0: No knowledge sources listed.
      • 1: Mentions knowledge sources, but the description is vague or the sources are not sufficiently authoritative.
      • 2: Clearly lists and explains the use of authoritative medical knowledge sources.
  3. 3.
    Medical Standards Alignment
    • Explanation: Benchmarks should follow internationally or industry-recognized medical standards (e.g., ICD, SNOMED CT, LOINC) when medical terminology is involved.
    • Justification: Adherence to medical standards ensures the clinical relevance and consistency, facilitating integration and comparison in reflect real-world medical practice.
    • Scoring:
      • 0: Does not mention any medical standards.
      • 1: Mentions the importance of standardization, but does not specify which standards are followed, or the description is unclear.
      • 2: Clearly states the adherence to recognized medical standards.
C.1.4 Evaluation Metrics and Dimensions
  1. 1.
    Validity of Core Metric
    • Explanation: The core performance metric should be clearly defined and closely related to the clinical task or medical capability being assessed.
    • Justification: Evaluation metrics directly shape the interpretation of results. Choosing suitable metric and explaining its relevance ensures a shared understanding and comprehensive interpretation.
    • Scoring:
      • 0: Core metrics are not clearly defined or have weak relevance to the task.
      • 1: Core metrics are clearly defined, but the choice and relevance of the metrics are not explained and justified.
      • 2: Core metrics are clearly defined, with justification for the relevance to the evaluation task.
  2. 2.
    Multi-dimensional Evaluation
    • Explanation: Apart from the correctness, benchmark developers should include evaluation of other important dimensions (e.g., safety, fluency).
    • Justification: In high-stakes medical domain, going beyond correctness is vital. Multi-dimensional evaluation offers a more comprehensive assessment of whether a model is truly reliable and trustworthy.
    • Scoring:
      • 0: Only consider the correctness of answers.
      • 1: Other evaluation dimensions are briefly mentioned, without specific evaluation methods.
      • 2: Multiple evaluation dimensions are clearly designed, with well defined evaluation methods.
  3. 3.
    Safety and Fairness Considerations
    • Explanation: Benchmark should include evaluation for potential risks (e.g., unsafe recommendation, toxicity) and bias (e.g., gender, ethnicity) in model outputs.
    • Justification: Evaluating risks and fairness facilitates bias-free, safe and equitable applications of LLMs in the medical domain, promoting responsible AI development and application.
    • Scoring:
      • 0: Does not consider safety and fairness evaluation.
      • 1: Mentions the importance of safety and fairness, but lacks specific evaluation methods.
      • 2: Designs dedicated assessment and test cases for safety and fairness.
C.2 Dataset Construction and Management
C.2.1 Data Source and Quality
  1. 1.
    Data source transparency and traceability
    • Explanation: Benchmark developers should clearly state the data source of the benchmark, along with relevant traceability information (e.g., data collection time frame, platforms).
    • Justification: Clear and traceable data sources are critical to ensure transparency and ethical data usage, which is especially important when sensitive data is involved. Also, it ensures reproducibility, enhancing the credibility of the benchmark.
    • Scoring:
      • 0: Does not state the original data source of the benchmark.
      • 1: States the data source of the benchmark, with incomplete traceability information.
      • 2: Clearly states the data source of the benchmark and provides detailed traceability information.
  2. 2.
    Data Source Reliability
    • Explanation: Benchmark developers should clearly describe the selection criteria and explain the reliability of the data source.
    • Justification: Collecting data from unreliable sources may lead to invalid results for medical applications. By explaining the reliability of the data source, credibility of the benchmark can be established.
    • Scoring:
      • 0: Does not describe the quality or selection criteria of the data source.
      • 1: Mentions the reliability of the data source, but does not explain the selection criteria.
      • 2: Clearly explains the selection criteria and describes the reliability of the data source.
  3. 3.
    Data Authenticity
    • Explanation: Benchmark developers should clearly specify whether the data comes from the real world scenarios, synthetically generated, or a mixture of both. For synthetically generated data, the construction process and verification for its authenticity (e.g., expert review) should also be described.
    • Justification: Real-world data reflects authentic scenarios but is harder to obtain, whereas synthetic data can be generated but requires careful verification. Clarifying data authenticity enhance transparency and ensures clinical relevance of the data used.
    • Scoring:
      • 0: Does not state the data source and explain its authenticity.
      • 1: States the data source, but lacks sufficient details on synthetic data generation or verification.
      • 2: Clearly states the data source, and provides a detailed description of the generation and verification process for synthetic data.
  4. 4.
    Dataset Representativeness
    • Explanation: The representativeness of key features (e.g., patient age, disease) of the dataset should be explained and statistically analyzed.
    • Justification: A benchmark that lacks representativeness may lead to bias in evaluation, reducing the clinical relevance, generalizability and fairness of the results.
    • Scoring:
      • 0: Does not mention the representativeness of the dataset.
      • 1: Provides qualitative descriptions of representativeness or partial and incomplete statistical analysis for some features.
      • 2: Provides quantitative statistical analysis of key features and compares them against the target population’s distribution.
  5. 5.
    Dataset Diversity
    • Explanation: The benchmark should have a diverse coverage, and provide quantitative evidence of its diversity of disease or medical departments covered.
    • Justification: Ensuring the dataset covers a variety helps comprehensively evaluate the model’s generalization ability, reducing bias in the evaluation results.
    • Scoring:
      • 0: Does not clearly define or explain the disease and medical department coverage.
      • 1: Clearly states the disease or medical department coverage, but coverage (defined in Equation 1) is less than average (21.8%).
      • 2: Clearly states the disease or medical department coverage with a coverage (defined in Equation 1) above average (21.8%).
C.2.2 Data Processing and Privacy Protection
  1. 1.
    Data Cleaning and Standardization
    • Explanation: The processes and steps of data preprocessing, including data cleaning and standardization, should be clearly described.
    • Justification: It ensures that the final dataset is well-structured, enhancing reliability. Also, it allows a better understanding of the construction process of the benchmark, ensuring transparency.
    • Scoring:
      • 0: Does not describe any data cleaning or standardization steps.
      • 1: The data preprocessing procedure is only briefly or partially mentioned.
      • 2: The data preprocessing procedure is clearly described, with details that cover all steps and aspects of the process.
  2. 2.
    Privacy Protection
    • Explanation: If sensitive data is used, it should be de-identified. Methods of de-identification should be described and compliance with relevant regulations (e.g., HIPAA) should be clearly stated.
    • Justification: Real-world clinical data may contain patient information. It is essential to ensure that the data and privacy protection aligns with ethical and legal standards, reducing the risk of privacy breaches and ensuring compliant data usage.
    • Scoring:
      • 0: No privacy protection measures are mentioned.
      • 1: Privacy protection measures are mentioned, but the description of methods or regulations is unclear.
      • 2: Sensitive data is not used. Otherwise, privacy protection measures are clearly described and compliance with relevant regulations is stated.
  3. 3.
    Data Format Clarity and Consistency
    • Explanation: The data in the dataset, including questions, cases, or task descriptions, should be written clearly and presented in a consistent format.
    • Justification: A clear and consistent format is essential for standardized evaluation. Inconsistent format may affect models’ interpretation, compromising the accuracy of the evaluation.
    • Scoring:
      • 0: Data format is disorganized and contains ambiguities.
      • 1: Minor inconsistencies or ambiguities exist in data format.
      • 2: Data format is clear and consistent.
  4. 4.
    Data Review and Audit
    • Explanation: The dataset construction process should include a review procedure with involvement of medical experts.
    • Justification: A dataset construction process without review mechanism is prone to errors. Careful review that involves medical experts can ensures clinical relevance, professionalism and reliable of the data.
    • Scoring:
      • 0: No review procedure is described.
      • 1: Briefly mentions the review procedure, or the review procedure is simple and without experts involvement.
      • 2: Provides a clear and detailed description of the review process, including the involvement of medical experts.
  5. 5.
    Quality of Reference Answer
    • Explanation: The benchmark should provide clear and accurate reference answers or scoring guidelines, and explain the construction and verification process (e.g., expert consensus).
    • Justification: Clear reference answers or scoring guidelines ensures transparent and accurate evaluation. By explaining how are reference answers or scoring guidelines formulated and verified, it also enhances the credibility of evaluation results.
    • Scoring:
      • 0: No reference answers or scoring guidelines are provided.
      • 1: Reference answers or scoring guidelines are provided, but the formulation or verification process is not explained.
      • 2: Provides clear reference answers or scoring guidelines, and explains its formulation and verification process in details.
  6. 6.
    Data Contamination Prevention
    • Explanation: Benchmark developers should take actions to identify, address and prevent potential data contamination issues of the data.
    • Justification: Data contamination may lead to inflated performance, which only reflect memorization instead of medical capability from the models. Preventing and controlling potential contamination ensures that the benchmark is effective, enhancing credibility and validity.
    • Scoring:
      • 0: Does not mention or consider potential data contamination.
      • 1: Mentions the risks of data contamination without concrete actions taken to handle data contamination.
      • 2: Clearly describes the procedures taken to prevent and control data contamination.
C.3 Technical Implementation and Evaluation Methodology
  1. 1.
    User-friendliness of evaluation tools
    • Explanation: Evaluation scripts or tools that is easy to obtain and use should be provided.
    • Justification: It ensures that users can use the benchmark conveniently, promoting benchmark adoption and ensuring fair, transparent, and consistent evaluation.
    • Scoring:
      • 0: No evaluation scripts or tools is provided.
      • 1: Provides evaluation scripts, but requires users to perform complex environment configurations or manual settings.
      • 2: Provides evaluation scripts or tools with detailed instructions for streamlined execution, that are user-friendly and require minimal user effort.
  2. 2.
    Technical Reproducibility
    • Explanation: Tools and technical documentation (e.g., environment configuration, dependency versions) should be provided.
    • Justification: The availability of evaluation tools and clear technical documentation ensures reproducibility of reported results, thereby enhancing the credibility of the benchmark.
    • Scoring:
      • 0: No information for technical reproducibility has been provided.
      • 1: Information for technical reproducibility is partially provided, but insufficient to guarantee reproducibility.
      • 2: Complete information for reproducibility is provided with detailed steps to replicate the results.
  3. 3.
    Provision of performance baselines
    • Explanation: Benchmark developers should provide multiple meaningful performance baselines, such as random, baseline models and human performance.
    • Justification: Providing different performance baselines allows comparison against the model’s performance, enabling a deeper understanding and better interpretability.
    • Scoring:
      • 0: No performance baselines are provided.
      • 1: Only one type of performance baseline is provided, or performance baselines are mentioned without specific data.
      • 2: At least two meaningful types of performance baselines are provided, with clear explanation of the measurement methods.
  4. 4.
    Reasoning Process Evaluation
    • Explanation: Apart from the final answer, the benchmark includes evaluations for the model’s reasoning process or explanation abilities.
    • Justification: In medical domain, understanding the model’s decision-making process is just as important as the final answer. Evaluating the reasoning process helps ensure that the model’s decisions are trustworthy and logically sound.
    • Scoring:
      • 0: Does not consider reasoning process evaluation.
      • 1: Mentions the importance of evaluating the reasoning process, but lacks specific evaluation methods.
      • 2: Designs assessment for reasoning process with clear evaluation methods and metrics.
  5. 5.
    Robustness Evaluation
    • Explanation: Benchmark should include testing for the model’s stability and robustness (e.g., input perturbations, adversarial samples)
    • Justification: In practical applications, models may encounter different variations of inputs. Robustness testing ensures model’s output is consistent and reliable under different conditions.
    • Scoring:
      • 0: Does not consider robustness evaluation.
      • 1: Mentions the importance of evaluating the robustness of models, but lacks specific evaluation
      • 2: Designs assessment for robustness with clear evaluation methods and metrics.
  6. 6.
    Generalization Capability Evaluation
    • Explanation: The benchmark design should help evaluate the generalization capability of models to unseen data or scenarios (e.g., careful train/test split, out-of-distribution testing).
    • Justification: Due to the high variability in medical scenarios, assessing a model’s generalization capability is essential for determining its reliability across different real-world scenarios, ensuring that the performance is not caused by overfitting.
    • Scoring:
      • 0: Does not consider generalization capability evaluation.
      • 1: Mentions the importance of evaluating the generalization capability of models, but lacks specific evaluation
      • 2: Designs mechanism for assessing the generalization capability.
  7. 7.
    Uncertainty Evaluation
    • Explanation: Benchmark should include evaluation for the model’s ability to recognize and express its own uncertainty (e.g., responding “I don’t know”)
    • Justification: Incorrect overconfident answers can be dangerous in high-stakes medical applications. A model’s ability to accurately identify and express uncertainty is critical for preventing incorrect and misleading decisions, enhancing system safety.
    • Scoring:
      • 0: Does not consider uncertainty evaluation.
      • 1: Mentions the importance of evaluating the uncertainty of models, but lacks specific evaluation
      • 2: Designs assessment for uncertainty with clear evaluation methods and metrics.
  8. 8.
    Evaluation Flexibility
    • Explanation: Evaluation code should have modular interface and support different models, including closed-source models that require API access and local open-source models.
    • Justification: It ensures that different types of models can be tested under the same interface.
    • Scoring:
      • 0: Supports only one type of model and is difficult to extend, requiring significant user effort to adapt.
      • 1: Supports only one type of model, but clear guidelines for extension is provided.
      • 2: Supports both closed-source and open-source models through a unified and user-friendly interface.
C.4 Benchmark Validity and Performance Verification
  1. 1.
    Knowledge and Skill Coverage
    • Explanation: Evidence (e.g., expert evaluation) is provided to demonstrate that the benchmark content sufficiently covers the medical knowledge and skills it claims to assess.
    • Justification: Sufficient coverage of the core medical competencies the benchmark aims to measure is the prerequisite for establishing content validity. When the content validity is supported with evidence, it enhances the rigor and trustworthiness of the evaluation.
    • Scoring:
      • 0: No evidence regarding coverage of knowledge and skills is provided.
      • 1: Provide brief content analysis or expert evaluation to demonstrate the coverage of knowledge and skills.
      • 2: Provides both comprehensive content analysis and expert evaluation to demonstrate the coverage of knowledge and skills.
  2. 2.
    Scenario Authenticity
    • Explanation: The evaluation tasks of the benchmark should effectively simulate and mirror real-world medical scenario.
    • Justification: Ensuring that the evaluation task closely mirrors the targeted clinical practice in real-world scenarios enhance the relevance of the benchmark. It helps assess whether the model can be transferred to real-world applications.
    • Scoring:
      • 0: The evaluation task have weak relevance to real clinical scenarios (e.g., abstract knowledge question answering)
      • 1: The evaluation task reflects a single clinical decision point, but does not constitute a complete workflow.
      • 2: The evaluation consists of multiple tasks that closely simulate realistic clinical workflows, with validation by domain experts.
  3. 3.
    Model Discrimination Ability
    • Explanation: Benchmark developers should provide experiment data and analysis, indicating that the benchmark can effectively distinguish the capability between different models.
    • Justification: An effective benchmark should be capable of differentiating and distinguishing models of varying capabilities. It provides meaningful insights into model strengths and weaknesses, guiding future improvements and development.
    • Scoring:
      • 0: No evidence is provided regarding the discrimination ability of the benchmark.
      • 1: The benchmark has been tested on some models and score differences are reported, but without statistical validation.
      • 2: The benchmark has been tested on several model, showing significant score differences with statistical validation.
  4. 4.
    Correlation with Clinical Performance
    • Explanation: Benchmark developers should provide evidence that preliminarily explores the correlation between benchmark performance and the model’s performance in actual clinical applications.
    • Justification: Validating whether benchmark results have meaningful indication on the model’s performance in real-world clinical scenarios ensures the external validity and practical value of the benchmark and results.
    • Scoring:
      • 0: Does not consider or mention the correlation between benchmark performance and clinical performance.
      • 1: Provides discussion about the correlation between benchmark performance and clinical performance theoretically without experiment data or concrete evidence.
      • 2: Preliminary research or evidence (e.g., small-scale clinical validation, consistency analysis between models and domain experts) is provided to explore and discuss the correlation between benchmark performance and clinical performance.
  5. 5.
    Internal Consistency
    • Explanation: For different sections or items of the benchmark evaluating the same capability, benchmark developers should demonstrate good internal consistency (e.g., Cronbach’s $\alpha$)
    • Justification: It ensures that different sections or items reliably assess the same capability, enabling meaningful comparisons and interpretation of results.
    • Scoring:
      • 0: Does not mention or conduct any internal consistency measurement.
      • 1: Mentions the importance of internal consistency, without concrete measurement.
      • 2: Conducts internal consistency measurement and demonstrates good consistency.
  6. 6.
    Statistical Significance Reporting
    • Explanation: When comparing the performance of different models, statistical significance (e.g., confidence intervals, and p-values) should be report.
    • Justification: Statistical significance testing allows for a more informed interpretation of results by differentiating the performance differences between models that stem from actual capabilities, and those arises from randomness and noise.
    • Scoring:
      • 0: No statistical significance is reported or considered.
      • 1: Reports simple statistics (e.g., mean and standard deviation of multiple executions) but does not conduct statistical testing.
      • 2: Reports and uses statistical significance tests (e.g., confidence interval).
C.5 Documentation, Openness and Governance
C.5.1 Documentation and Transparency
  1. 1.
    Documentation Completeness
    • Explanation: Benchmark developers should provide a clear and comprehensive documentation of the benchmark, systematically describing the relevant details, including the design, objectives, scope, construction process, task definitions and evaluation procedures.
    • Justification: A complete and clear documentation help users understand and use the benchmark properly, enhancing the usability, reproducibility and transparency.
    • Scoring:
      • 0: The documentation is missing or overly simple.
      • 1: The documentation exists, but some important details are missing or unclearly described.
      • 2: The documentation is comprehensive and clear, covering all key components.
  2. 2.
    Clarity of Evaluation Guidelines
    • Explanation: Benchmark developers should provide evaluation guidelines, with definitions of evaluation metrics, detailed scoring criteria, and easy-to-follow usage instructions for users to replicate the evaluation process.
    • Justification: Clear evaluation guidelines help users better understand how model performance is quantified, ensuring a shared interpretation and understanding. It also enhances the consistency by documenting how to replicate the evaluation process.
    • Scoring:
      • 0: No evaluation guidelines are provided.
      • 1: Provides evaluation guidelines, but some areas are only briefly described.
      • 2: Provides a clear and comprehensive evaluation guidelines.
  3. 3.
    Discussion of Limitations and Risks
    • Explanation: The benchmark developers should openly discuss the limitations and potential social risks of the benchmark.
    • Justification: Disclosing limitations and risks demonstrates scientific rigor and responsibility. This transparency helps prevent users from misunderstanding and misusing the benchmark, especially in high-stakes contexts.
    • Scoring:
      • 0: Discussion of limitations and risks is not provided.
      • 1: Limitations and risks are briefly mentioned without in-depth discussion.
      • 2: In-depth and comprehensive discussion of limitations and risks is provided.
  4. 4.
    Peer Review
    • Explanation: The benchmark and its corresponding paper was accepted at peer-reviewed venue.
    • Justification: Going through peer review process means that the design, validity and results of a benchmark has been rigorous evaluated. It enhances credibility and ensures quality.
    • Scoring:
      • 0: The benchmark has not been accepted at a peer-reviewed venue.
      • 1: The benchmark is under review or published on platform without strict peer review (e.g., arXiv preprints).
      • 2: The benchmark has been published at a peer-reviewed venue.
C.5.2 Openness and Accessibility
  1. 1.
    Accessibility of Evaluation Code and Data
    • Explanation: Access to the evaluation code and data, which can be shared within legal and ethical boundaries, is provided (e.g. on platforms like GitHub or Hugging Face) along with the applicable license.
    • Justification: Accessible code and data are prerequisites for reproducibility. Moreover, it allows the community to review, improve, and expand the benchmarks.
    • Scoring:
      • 0: Access to the evaluation code and data is not provided.
      • 1: The evaluation code and data are partially accessible, or accessible without clear licensing.
      • 2: The evaluation code and data are fully accessible and the applicable license is clearly specified.
  2. 2.
    Usage and Citation Guidelines
    • Explanation: Clear guidelines are provided to standardize benchmark usage, result reporting, and correct citation formats in academic papers or technical reports.
    • Justification: Proper usage and citation guidelines help maintain academic integrity. It also promote standardized reporting of results, facilitating comparisons in further research.
    • Scoring:
      • 0: No usage or citation guidelines are provided.
      • 1: Partial guidelines are provided but are incomplete.
      • 2: Clear and complete usage and citation guidelines are provided.
C.5.3 Continuous Maintenance and Governance
  1. 1.
    Update and Version Management
    • Explanation: A feedback channel should be maintained for users to report problems, provide feedback and suggestions.
    • Justification: Maintaining an effective feedback channel allows users to provide feedback when issues with the benchmark are discovered. This is crucial for continuously improving benchmark quality and fixing potential bugs and issues.
    • Scoring:
      • 0: No update plans or version management.
      • 1: Intention to update is mentioned, without any concrete actions.
      • 2: There is a clear plan for update and version management.
  2. 2.
    Feedback Channel for Users
    • Explanation: A feedback channel should be maintained for users to report problems, provide feedback and suggestions.
    • Justification: Maintaining an effective feedback channel allows users to provide feedback when issues with the benchmark are discovered. This is crucial for continuously improving benchmark quality and fixing potential bugs and issues.
    • Scoring:
      • 0: No public feedback channel is established.
      • 1: A public feedback channel is set up, but responses are either delayed or absent.
      • 2: A public feedback channel is set up with promptly response to user feedback.
  3. 3.
    Long-term Maintenance Responsibility
    • Explanation: The individuals, teams, or institutions responsible for its long-term maintenance and development should be clearly stated.
    • Justification: Clarifying who holds long-term responsibility reassures the community that it will be actively supported and improved, ensuring usability and credibility.
    • Scoring:
      • 0: No long-term maintenance responsibility is mentioned.
      • 1: The responsible party is implied, but not explicitly mentioned.
      • 2: Details of the person responsible for long-term maintenance are clearly stated.